# A Galactic Transformation—Understanding the SMC’s Structural and Kinematic Disequilibrium

### v1

@@ 3.1 The Simulated SMC’s Stellar Structure
To quantify the tidal extension, we compute the SMC’s spherically averaged stellar volume density profiles at different simulation epochs (bottom right panel of Figure 5). At MW Infall, the SMC’s stellar density profile is well described by an exponential disk. However, post-collision, the stellar density profile is instead a combination of at least two power laws. The SMC’s stellar density at $R\lesssim 3$ kpc decreases by a factor of 2-3. Correspondingly, the stellar densities in the outer regions, $R\gtrsim 3$ kpc, increase by a factor of 2-3. The power-law break radius of the post-collision profiles is between 242-4 kpc, depending on the time elapsed since the collision. The small break radii indicate that the internal structure of the SMC is significantly affected by the LMC’s tides. Indeed, there is observational evidence that the LMC’s tides are significantly influencing the SMC’s stellar distribution at $R<4$ kpc (Z21), which we discuss in more detail in section 3.2.
As mentioned in section 2.1, we utilize the B12 Model 1 SMC as a control, where the SMC is still subjected to MW and LMC tides, but does not collide with the LMC. Here, the Model 2 MW Infall stellar density profile is representative of the SMC profile in Model 1 at present day. Hence, neither weak LMC tides nor MW tides can significantly affect the SMC’s stellar density profile at $R<4$ kpc. A collision is required to significantly disrupt the internal stellar morphology of the SMC.
@@
3.1.1 The SMC’s LoS Elongation
Figure 6: The simulated SMC’s principal axis ratios computed at varying radii as a function of time. Top row shows the major-to-minor axis ratio (a/c) and the bottom row shows the intermediate-to-minor axis ratio. Columns show the results for Model 2 (left) and the Model 1 control (right). For Model 2, different epochs (SMC-LMC collision, 100 Myr and 200 Myr post-collision) are marked by dotted vertical lines. Post-collision, for $R<5$ kpc the SMC stellar structure is a triaxial ellipsoid. At larger spatial scales ($R<15$ kpc), $a:c$ becomes as high as $5.5:1$, indicating that an axis has become significantly more elongated, consistent with observations, which find $a:c>4:1$ (Crowl2001). Without a collision (Model 1, right), axis ratios remain <2.5:1<2.5:1 for all spatial scales, which is inconsistent with observations.
Measuring the SMC’s LoS depth is challenging in both observations and simulations. With observations, the inferred values depend significantly on the criteria adopted for quantifying the LoS extent, photometric depth of the survey, as well as the dust extinction prescription and the stellar population used (see references in section 1). With simulations, it is challenging to define a LoS because the observed SMC’s on-sky inclination is not well constrained.
Given the aforementioned hurdles, we choose to study the simulated SMC’s LoS extent by computing the axis ratios of the stellar distribution, since the axis ratios are independent of the assumed orientation. Several observational studies (Crowl2001; Subramanian2012) have measured the axis ratios of the SMC’s stellar distribution, finding the major-to-minor axis ratio ($a:c$) to be large (>4:1>4:1), which is consistent with a large LoS depth.
A common way to compute the axis ratios of a stellar distribution is by constructing the moment of inertia tensor ($\mathbf{\tilde{I}}$) (Paz2006; Pejcha2009; Subramanian2012). We center the simulated SMC at the stellar density center and, following Subramanian2012, we write the cartesian representation of ($\mathbf{\tilde{I}}$) as:
@@ 3.1.1 The SMC’s LoS Elongation
The axis ratios at all times in Model 1 and prior to the collision in Model 2 are (22.5):(1.52):1(2-2.5):(1.5-2):1. This indicates that the SMC has undergone significant vertical heating caused by previous distant (>10>10 kpc separation at pericenter) encounters between the Clouds that have already happened before they infall into the MW (see section 3.2 for more details).
In Model 2, post-collision, axis ratios are similar to the pre-collision values for $R<5$ kpc, indicating a triaxial main body. This finding is qualitatively consistent with observations; for e.g. Subramanian2012 used distances to SMC’s RR Lyrae stars in the SMC from the OGLE III survey (Soszynski2010) and concluded that the stellar distribution is triaxial, but the exact axis ratios are sensitive to the size of the spatial region used to select the stars.
However, for Model 2 at large spatial scales ($R<15$ kpc), the major-to-minor axis ratio achieves values as high as $5.5:1$. This indicates that the post-collision SMC has developed a long axis at large radii that was non-existent pre-collision. The simulated post-collision axis ratios at $R<15$ kpc are qualitatively consistent with observations; e.g. Crowl2001 used distances to SMC’s star clusters and measured axis ratios of $1:2:4$ for $R<12$ kpc. In Model 1, the major to minor axis ratio remains smaller than $2.5:1$ at all times and for all spatial scales, which is discrepant with the observed SMC’s LoS depth.
@@
3.1.2 Orientation of the SMC’s Major Axis
Figure 7: The left panel shows a projection of the Model 2 simulated SMC’s stellar density distribution 200 Myr after the SMC-LMC collision. The x’ and y’ axis correspond to the SMC plane of rotation. The stellar density centers of the SMC and LMC are marked by the red star and green circle, respectively. The LMC-SMC bridge and the SMC’s tidal tail can be clearly seen. The latter is likely responsible for the SMC’s large LoS depth (DB2012). The right panel shows a cartoon illustration (not to scale) of how we might be observing the SMC. Our LoS likely passes through the SMC’s tidal tail, leading to a large observed LoS depth. The LMC-SMC bridge, being misaligned with the tidal tail, does not significantly affect the LoS depth.
@@ 3.1.2 Orientation of the SMC’s Major Axis
Because the long axis appears only at large spatial scales, it likely corresponds to the SMC’s tidal structures born out of the SMC-LMC collision. In Figure 7 left panel, we show the Model-2 SMC’s stellar surface density distribution 200 Myr post-collision, in the SMC’s plane of rotation. Two tidal structures can be clearly seen: a bridge of stars connecting the SMC and LMC; and a tidal tail of stars along the positive x’-axis. These are classic examples of tidal bridges and tails that form out of binary galaxy interactions (Toomre1972).
@@
Previous theoretical studies (Yoshizawa2003; Connors2006; DB2012) have also suggested that the SMC’s tidal tail is the origin of the large LoS depth; this tidal tail is referred to as the “counter bridge” in DB2012. However, the aforementioned studies did not consider a direct collision between the Clouds, and instead considered more distant encounters (closest separation $\gtrsim 6$ kpc). In this work, we have shown that a direct SMC-LMC collision also leads to a similar tidal tail, thereby explaining the large axis ratios (or equivalently a large LoS depth).
The viewing perspective suggested in Figure 7 requires the bridge to be misaligned with the tidal tail. Observationally, the angle between the bridge and the LoS ($\theta_{\rm obs}$) is computed using the Galactocentric coordinates of the SMC and the Sun (Kallivayalil2013):
@@
Note that the tidal tail discussed in the context of SMC’s LoS depth is different from the MS and its stellar counterpart (Chandra2023; Zaritsky2025). A tidal stream analogous to the MS forms in the B12 simulation at previous SMC-LMC pericentric passages (see also Yoshizawa2003; Connors2006; DB2012; Pardy2018). Whereas, the tidal tail we discuss here is formed out of the recent SMC-LMC collision, <200<200 Myr ago.
@@
3.2 The Simulated SMC’s Stellar Kinematics
@@
Z21 analyzed Gaia DR2 (Brown2018) PMs of the SMC’s red giant stars combined with LoS velocities from Dobbie2014 and the APOGEE survey (Majewski2017). They found that the internal rotation field of the SMC is very complex. Stars in the inner $\sim 1$ kpc show small rotation ($\approx 10$ km s-1), whereas the kinematics beyond $\sim 1$ kpc are dominated by radially outward motions (see also Niederhofer2021), likely due to the LMC’s tides. This suggests that the LMC’s tides must be dominant over the SMC’s own gravitational force well inside the SMC’s main body (R <4<4 kpc).
@@
In Figure 9, the simulated Model 2 SMC’s stellar velocity field is visualized in the plane of rotation at different epochs. Pre-collision (top left panel), the SMC has a well defined stellar rotation field. Post-collision (top right and bottom left panels), the stellar rotation field is significantly disrupted. Most of the stars at R >2>2 kpc are moving radially outwards. Further, the direction of maximal outward motion is consistent with $\boldsymbol{R_{\rm LMC-SMC}}$, indicating that the motions are driven by the LMC’s tides. There is some remnant rotation with a small amplitude (<10<10 km s-1) in the inner $\sim 1$ kpc. This remnant rotation suggests that stars in the inner 121-2 kpc are bound to the SMC. The simulated post-collision SMC kinematics are consistent with the remnant rotation and the tidal expansion inferred by Z21. Hence, the SMC-LMC collision can explain the observed SMC’s stellar kinematics.
@@
Observational studies (e.g. Harris2006, hereafter HZ06) have found that the SMC is a highly dispersion dominated galaxy, where the ratio of the peak rotation velocity to the total velocity dispersion ($\frac{v_{\phi,\rm{max}}}{\sigma}$) is significantly less than 0.6. Thus, not only are the SMC’s stars moving radially outwards (Z21), their kinematics are also significantly heated. Next, we investigate whether the SMC-LMC collision can explain the heating of the SMC’s stellar kinematics.
@@
It has been shown that multiple distant encounters can significantly heat a disk (e.g. Mayer2001; Lokas2015; Semczuk2018). However, distant encounters with the MW and LMC (Model 1) or just the LMC (Model 2, MW Infall) are not sufficient to make the SMC dispersion dominated at the level seen in observations.
@@ 3.3 The Simulated SMC’s Gas Kinematics
Observationally, gas 3D kinematics are unknown, leaving only the LoS velocities to study the internal gas kinematics of the SMC. The observed SMC shows a HI LoS velocity gradient of 6010060-100 km s-1 across an on-sky extent of $\approx 8^{\circ}$ (DiTeodoro2019). Rotating disk models (e.g. Marel2001; Marel2002) have been fit to the LoS velocity field, and a rotation curve is inferred from the data, with a peak amplitude of $\approx 50$ km s-1 (DiTeodoro2019).
@@
Pre-collision, we found that the stellar kinematic center, stellar systemic velocity, and the stellar disk angular momentum vector are adequate representations of the gas kinematics in both Model 2 and Model 1. However, post-collision in Model 2, we were unable to reliably determine the SMC gas kinematic center and the gas angular momentum vector. Despite our best efforts at fine tuning the iterative algorithms described in section 2.2, we failed to converge. Moreover, the inferred gas kinematic center and the gas angular momentum vector are very sensitive to the choice of hyper-parameters in the iterative algorithms. Being unsuccessful at centering and aligning the gas distribution already indicates that the gas velocity field is significantly more disturbed than the stellar velocity field. To analyze the SMC’s gas kinematics, we choose the stellar kinematic center to center the gas spatial distribution post-collision. We have checked that the choice of centering does not affect our final conclusions for this section.
@@
Since the post-collision gas kinematics are dominated by radially outward motions, viewing such radial motions from an inclined perspective can result in a LoS velocity gradient that could be confused as a rotation signature (see SylosLabini2019). Next, we investigate whether the post-collision simulated SMC’s gas shows LoS velocity gradients at a particular viewing perspective.
We assume a viewing perspective where the longest stellar axis (section 3.1) is a proxy for the LoS. Such a viewing perspective is consistent with the observed SMC’s large LoS depth. We use the transformation matrix $\mathbf{T}$ (eq. 5) to accordingly rotate the gas distribution and kinematics. The new coordinate axes are called $(x^{\prime\prime},y^{\prime\prime},z^{\prime\prime})$, where $z^{\prime\prime}$ is along the longest stellar axis and serves as a proxy for the LoS.
@@
Contours of both stellar and gas densities are shown in Figure 12. At MW Infall, the gas density contours are significantly more extended as compared to the stellar density contours, largely due to the more extended gas disk adopted in the initial conditions (see section 2.1). Pre-collision, the gas zero velocity curve (defined by the locus of points in the $x^{\prime\prime}-y^{\prime\prime}$ plane that have an LoS velocity equal to the systemic velocity) is well centered with both the stellar and gas density contours. However, post-collision, the zero velocity curve is displaced from the innermost stellar and gas density contour by $\sim 1$ kpc (see the Collision + 200 Myr epoch in particular). Hence, the SMC-LMC collision can explain the separation between the observed SMC’s gas LoS zero velocity curve and its stellar photometric center (see Figure 1, right panel).
@@
If the observed SMC’s gas kinematics does not have a significant rotation, then it is not reasonable to use the peak of the gas velocity gradient as a rotation signature in the BTFR. This would resolve the discrepancy between the SMC’s gas and stellar kinematics. We conclude that the SMC does not reside on the BTFR, irrespective of whether stellar or gas kinematics is considered. The SMC is a galaxy with a very high degree of kinematic disequilibrium.
In this section, we have seen that post SMC-LMC collision, the simulated SMC’s gas is significantly more disturbed as compared to the SMC’s stars. Clearly, there must be forces apart from the LMC’s tidal forces that preferentially act on the SMC’s gas and not stars. These forces are hydrodynamic forces, which are a consequence of the collisional nature of the galaxies’ gas components. The role of hydrodynamics during the collision is discussed in section 4.1.
4 Discussion
We have shown using the B12 hydrodynamic simulations that the recent SMC-LMC collision can explain observations of the SMC’s large LoS depth, lack of stellar rotation, and discrepant stellar and gas centers. Post-collision, the simulated SMC’s gas kinematics has significant radially outward motions, which can manifest as a LoS velocity gradient if viewed at an inclined perspective. Such a gradient can be misinterpreted as a rotating disk. Consequently, the SMC does not reside on the BTFR, and is in a very high state of disequilibrium. Next, we discuss the following additional aspects of the SMC’s disequilibrium: (i) effect of LMC’s ram pressure on the SMC’s gas (section 4.1); (ii) The distribution of SMC’s gas along the LoS (section 4.2); and (iii) implications for measuring the SMC’s DM content (section 4.3). We also compare our analyses and results with previous works in section 4.4. Finally, we present the limitations of our work and discuss future prospects (section 4.5).
4.1 Impact of LMC Ram Pressure on the SMC’s Gas
Figure 13: Ram pressure (green solid line) exerted on the SMC’s gas disk by the LMC’s gas disk during an impulsive SMC-LMC collision. The orange dashed line denotes the restoring thrust applied by the SMC’s mass distribution on its gas. The ram pressure from the LMC is more than an order of magnitude larger than the SMC’s restoring thrust. Due to ram pressure, the SMC’s gas gets an impulsive velocity kick of $\approx 30$ km s-1. Hence, ram pressure can potentially explain the lack of remnant rotation in gas (Figure 11), and the spatial offset between the SMC’s stellar density center and the gas LoS zero velocity curve (see Figure 1 right panel and Figure 12). The ram pressure exerted by the LMC’s ISM on the SMC during the SMC-LMC collision is much larger than the typical ram pressure values in galaxy CGMs (Emerick2016) and on the upper end of the ram pressure exerted by ICMs (Xie2025).

### v2

@@ III.1 The Simulated SMC’s Stellar Structure
To quantify the tidal extension, we compute the SMC’s spherically averaged stellar volume density profiles at different simulation epochs (bottom right panel of Figure 5). At MW Infall, the SMC’s stellar density profile is well described by an exponential disk. However, post-collision, the stellar density profile is instead a combination of at least two power laws. The SMC’s stellar density at $R\lesssim 3$ kpc decreases by a factor of 2-3. Correspondingly, the stellar densities in the outer regions, $R\gtrsim 3$ kpc, increase by a factor of 2-3. The power-law break radius of the post-collision profiles is between 242-4 kpc, depending on the time elapsed since the collision. The small break radii indicate that the internal structure of the SMC is significantly affected by the LMC’s tides. Indeed, there is observational evidence that the LMC’s tides are significantly influencing the SMC’s stellar distribution at $R<4$ kpc (Z21), which we discuss in more detail in section III.2.
As mentioned in section II.1, we utilize the B12 Model 1 SMC as a control, where the SMC is still subjected to MW and LMC tides, but does not collide with the LMC. Here, the Model 2 MW Infall stellar density profile is representative of the SMC profile in Model 1 at present day. Hence, neither weak LMC tides nor MW tides can significantly affect the SMC’s stellar density profile at $R<4$ kpc. A collision is required to significantly disrupt the internal stellar morphology of the SMC.
@@
III.1.1 The SMC’s LoS Elongation
Figure 6: The simulated SMC’s principal axis ratios computed at varying radii as a function of time. Top row shows the major-to-minor axis ratio (a/c) and the bottom row shows the intermediate-to-minor axis ratio. Columns show the results for Model 2 (left) and the Model 1 control (right). For Model 2, different epochs (SMC-LMC collision, 100 Myr and 200 Myr post-collision) are marked by dotted vertical lines. Post-collision, for $R<5$ kpc the SMC stellar structure is a triaxial ellipsoid. At larger spatial scales ($R<15$ kpc), $a:c$ becomes as high as $5.5:1$, indicating that an axis has become significantly more elongated, consistent with observations, which find $a:c>4:1$ (Crowl et al., 2001). Without a collision (Model 1, right), axis ratios remain <2.5:1<2.5:1 for all spatial scales, which is inconsistent with observations.
Measuring the SMC’s LoS depth is challenging in both observations and simulations. With observations, the inferred values depend significantly on the criteria adopted for quantifying the LoS extent, photometric depth of the survey, as well as the dust extinction prescription and the stellar population used (see references in section I). With simulations, it is challenging to define a LoS because the observed SMC’s on-sky inclination is not well constrained.
Given the aforementioned hurdles, we choose to study the simulated SMC’s LoS extent by computing the axis ratios of the stellar distribution, since the axis ratios are independent of the assumed orientation. Several observational studies (Crowl et al., 2001; Subramanian and Subramaniam, 2012) have measured the axis ratios of the SMC’s stellar distribution, finding the major-to-minor axis ratio ($a:c$) to be large (>4:1>4:1), which is consistent with a large LoS depth.
A common way to compute the axis ratios of a stellar distribution is by constructing the moment of inertia tensor ($\mathbf{\tilde{I}}$) (Paz et al., 2006; Pejcha and Stanek, 2009; Subramanian and Subramaniam, 2012). We center the simulated SMC at the stellar density center and, following Subramanian and Subramaniam (2012), we write the cartesian representation of ($\mathbf{\tilde{I}}$) as:
@@ III.1.1 The SMC’s LoS Elongation
The axis ratios at all times in Model 1 and prior to the collision in Model 2 are (22.5):(1.52):1(2-2.5):(1.5-2):1. This indicates that the SMC has undergone significant vertical heating caused by previous distant (>10>10 kpc separation at pericenter) encounters between the Clouds that have already happened before they infall into the MW (see section III.2 for more details).
In Model 2, post-collision, axis ratios are similar to the pre-collision values for $R<5$ kpc, indicating a triaxial main body. This finding is qualitatively consistent with observations; for e.g. Subramanian and Subramaniam (2012) used distances to SMC’s RR Lyrae stars in the SMC from the OGLE III survey (Soszyński et al., 2010) and concluded that the stellar distribution is triaxial, but the exact axis ratios are sensitive to the size of the spatial region used to select the stars.
However, for Model 2 at large spatial scales ($R<15$ kpc), the major-to-minor axis ratio achieves values as high as $5.5:1$. This indicates that the post-collision SMC has developed a long axis at large radii that was non-existent pre-collision. The simulated post-collision axis ratios at $R<15$ kpc are qualitatively consistent with observations; e.g. Crowl et al. (2001) used distances to SMC’s star clusters and measured axis ratios of $1:2:4$ for $R<12$ kpc. In Model 1, the major to minor axis ratio remains smaller than $2.5:1$ at all times and for all spatial scales, which is discrepant with the observed SMC’s LoS depth.
@@
III.1.2 Orientation of the SMC’s Major Axis
Figure 7: The left panel shows a projection of the Model 2 simulated SMC’s stellar density distribution 200 Myr after the SMC-LMC collision. The x’ and y’ axis correspond to the SMC plane of rotation. The stellar density centers of the SMC and LMC are marked by the red star and green circle, respectively. The LMC-SMC bridge and the SMC’s tidal tail can be clearly seen. The latter is likely responsible for the SMC’s large LoS depth (Diaz and Bekki, 2012). The right panel shows a cartoon illustration (not to scale) of how we might be observing the SMC. Our LoS likely passes through the SMC’s tidal tail, leading to a large observed LoS depth. The LMC-SMC bridge, being misaligned with the tidal tail, does not significantly affect the LoS depth.
@@ III.1.2 Orientation of the SMC’s Major Axis
Because the long axis appears only at large spatial scales, it likely corresponds to the SMC’s tidal structures born out of the SMC-LMC collision. In Figure 7 left panel, we show the Model-2 SMC’s stellar surface density distribution 200 Myr post-collision, in the SMC’s plane of rotation. Two tidal structures can be clearly seen: a bridge of stars connecting the SMC and LMC; and a tidal tail of stars along the positive x’-axis. These are classic examples of tidal bridges and tails that form out of binary galaxy interactions (Toomre and Toomre, 1972).
@@
Previous theoretical studies (Yoshizawa and Noguchi, 2003; Connors et al., 2006; Diaz and Bekki, 2012) have also suggested that the SMC’s tidal tail is the origin of the large LoS depth; this tidal tail is referred to as the “counter bridge” in Diaz and Bekki (2012). However, the aforementioned studies did not consider a direct collision between the Clouds, and instead considered more distant encounters (closest separation $\gtrsim 6$ kpc). In this work, we have shown that a direct SMC-LMC collision also leads to a similar tidal tail, thereby explaining the large axis ratios (or equivalently a large LoS depth).
The viewing perspective suggested in Figure 7 requires the bridge to be misaligned with the tidal tail. Observationally, the angle between the bridge and the LoS ($\theta_{\rm obs}$) is computed using the Galactocentric coordinates of the SMC and the Sun (Kallivayalil et al., 2013):
@@
Note that the tidal tail discussed in the context of SMC’s LoS depth is different from the MS and its stellar counterpart (Chandra et al., 2023; Zaritsky et al., 2025). A tidal stream analogous to the MS forms in the B12 simulation at previous SMC-LMC pericentric passages (see also Yoshizawa and Noguchi, 2003; Connors et al., 2006; Diaz and Bekki, 2012; Pardy et al., 2018). Whereas, the tidal tail we discuss here is formed out of the recent SMC-LMC collision, <200<200 Myr ago.
@@
III.2 The Simulated SMC’s Stellar Kinematics
@@
Z21 analyzed Gaia DR2 (Gaia Collaboration et al., 2018) PMs of the SMC’s red giant stars combined with LoS velocities from Dobbie et al. (2014) and the APOGEE survey (Majewski et al., 2017). They found that the internal rotation field of the SMC is very complex. Stars in the inner $\sim 1$ kpc show small rotation ($\approx 10$ km s-1), whereas the kinematics beyond $\sim 1$ kpc are dominated by radially outward motions (see also Niederhofer et al., 2021; Dias et al., 2021, 2022; Parisi et al., 2024), likely due to the LMC’s tides. This suggests that the LMC’s tides must be dominant over the SMC’s own gravitational force well inside the SMC’s main body (R <4<4 kpc). De Leo et al. (2020) also arrived at a similar conclusion after analyzing the PMs and LoS velocities of a sample of 3000 SMC red giant stars.
@@
In Figure 9, the simulated Model 2 SMC’s stellar velocity field is visualized in the plane of rotation at different epochs. Pre-collision (top left panel), the SMC has a well defined stellar rotation field. Post-collision (top right and bottom left panels), the stellar rotation field is significantly disrupted. Most of the stars at R >2>2 kpc are moving radially outwards. Further, the direction of maximal outward motion is consistent with $\boldsymbol{R_{\rm LMC-SMC}}$, indicating that the motions are driven by the LMC’s tides. There is some remnant rotation with a small amplitude (<10<10 km s-1) in the inner $\sim 1$ kpc. This remnant rotation suggests that stars in the inner 121-2 kpc are bound to the SMC. At 100 Myr post collision, the simulated SMC’s total mass enclosed within a radius of 2 kpc is $6.8\times 10^{8}$ M. A rough estimate of the circular velocity at 2 kpc can be obtained by assuming spherical symmetry of the mass enclosed. This circular velocity turns out to be $\approx 38$ km s-1, reaffirming that stars at $R\lesssim 2$ kpc are likely bound to the SMC.
The simulated post-collision SMC kinematics are consistent with the remnant rotation and the tidal expansion inferred by Z21. Hence, the SMC-LMC collision can explain the observed SMC’s stellar kinematics.
@@
Observational studies (e.g. Harris and Zaritsky, 2006, hereafter HZ06) have found that the SMC is a highly dispersion dominated galaxy, where the ratio of the peak rotation velocity to the total velocity dispersion ($\frac{v_{\phi,\rm{max}}}{\sigma}$) is significantly less than 0.6. Thus, not only are the SMC’s stars moving radially outwards (Z21), their kinematics are also significantly heated. Next, we investigate whether the SMC-LMC collision can explain the heating of the SMC’s stellar kinematics.
@@
It has been shown that multiple distant encounters can significantly heat a disk (e.g. Mayer et al., 2001; Kazantzidis et al., 2011, 2013, 2017; Lokas, 2012; Łokas et al., 2015; Semczuk et al., 2018). However, distant encounters with the MW and LMC (Model 1) or just the LMC (Model 2, MW Infall) are not sufficient to make the SMC dispersion dominated at the level seen in observations.
@@ III.3 The Simulated SMC’s Gas Kinematics
Observationally, gas 3D kinematics are unknown, leaving only the LoS velocities to study the internal gas kinematics of the SMC. The observed SMC shows a HI LoS velocity gradient of 6010060-100 km s-1 across an on-sky extent of $\approx 8^{\circ}$ (Di Teodoro et al., 2019). Rotating disk models (e.g. van der Marel and Cioni, 2001; van der Marel et al., 2002) have been fit to the LoS velocity field, and a rotation curve is inferred from the data, with a peak amplitude of $\approx 50$ km s-1 (Di Teodoro et al., 2019).
@@
Pre-collision, we found that the stellar kinematic center, stellar systemic velocity, and the stellar disk angular momentum vector are adequate representations of the gas kinematics in both Model 2 and Model 1. However, post-collision in Model 2, we were unable to reliably determine the SMC gas kinematic center and the gas angular momentum vector. Despite our best efforts at fine tuning the iterative algorithms described in section II.2, we failed to converge. Moreover, the inferred gas kinematic center and the gas angular momentum vector are very sensitive to the choice of hyper-parameters in the iterative algorithms. Being unsuccessful at centering and aligning the gas distribution already indicates that the gas velocity field is significantly more disturbed than the stellar velocity field. To analyze the SMC’s gas kinematics, we choose the stellar kinematic center to center the gas spatial distribution post-collision. We have checked that the choice of centering does not affect our final conclusions for this section.
@@
Since the post-collision gas kinematics are dominated by radially outward motions, viewing such radial motions from an inclined perspective can result in a LoS velocity gradient that could be confused as a rotation signature (see Sylos Labini et al., 2019). Next, we investigate whether the post-collision simulated SMC’s gas shows LoS velocity gradients at a particular viewing perspective.
We assume a viewing perspective where the longest stellar axis (section III.1) is a proxy for the LoS. Such a viewing perspective is consistent with the observed SMC’s large LoS depth. We use the transformation matrix $\mathbf{T}$ (eq. 5) to accordingly rotate the gas distribution and kinematics. The new coordinate axes are called $(x^{\prime\prime},y^{\prime\prime},z^{\prime\prime})$, where $z^{\prime\prime}$ is along the longest stellar axis and serves as a proxy for the LoS.
@@
Contours of both stellar and gas densities are shown in Figure 12. At MW Infall, the gas density contours are significantly more extended as compared to the stellar density contours, largely due to the more extended gas disk adopted in the initial conditions (see section II.1). Pre-collision, the gas zero velocity curve (defined by the locus of points in the $x^{\prime\prime}-y^{\prime\prime}$ plane that have an LoS velocity equal to the systemic velocity) is well centered with both the stellar and gas density contours. However, post-collision, the zero velocity curve is displaced from the innermost stellar and gas density contour by $\sim 1$ kpc (see the Collision + 200 Myr epoch in particular). Hence, the SMC-LMC collision can explain the separation between the observed SMC’s gas LoS zero velocity curve and its stellar photometric center (see Figure 1, right panel).
@@
If the observed SMC’s gas kinematics does not have a significant rotation, then it is not reasonable to use the peak of the gas velocity gradient as a rotation signature in the BTFR. This would resolve the discrepancy between the SMC’s gas and stellar kinematics. We conclude that the SMC does not reside on the BTFR, irrespective of whether stellar or gas kinematics is considered. The SMC is a galaxy with a very high degree of kinematic disequilibrium (see also De Leo et al. 2020).
In this section, we have seen that post SMC-LMC collision, the simulated SMC’s gas is significantly more disturbed as compared to the SMC’s stars. Clearly, there must be forces apart from the LMC’s tidal forces that preferentially act on the SMC’s gas and not stars. These forces are hydrodynamic forces, which are a consequence of the collisional nature of the galaxies’ gas components. The role of hydrodynamics during the collision is discussed in section IV.1.
IV Discussion
We have shown using the B12 hydrodynamic simulations that the recent SMC-LMC collision can explain observations of the SMC’s large LoS depth, lack of stellar rotation, and discrepant stellar and gas centers. Post-collision, the simulated SMC’s gas kinematics has significant radially outward motions, which can manifest as a LoS velocity gradient if viewed at an inclined perspective. Such a gradient can be misinterpreted as a rotating disk. Consequently, the SMC does not reside on the BTFR, and is in a very high state of disequilibrium. Next, we discuss the following additional aspects of the SMC’s disequilibrium: (i) effect of LMC’s ram pressure on the SMC’s gas (section IV.1); (ii) The distribution of SMC’s gas along the LoS (section IV.2); and (iii) implications for measuring the SMC’s DM content (section IV.3). We also compare our analyses and results with previous works in section IV.4. Finally, we present the limitations of our work and discuss future prospects (section IV.5).
IV.1 Impact of LMC Ram Pressure on the SMC’s Gas
Figure 13: Ram pressure (green solid line) exerted on the SMC’s gas disk by the LMC’s gas disk during an impulsive SMC-LMC collision. The orange dashed line denotes the restoring thrust applied by the SMC’s mass distribution on its gas. The ram pressure from the LMC is more than an order of magnitude larger than the SMC’s restoring thrust. Due to ram pressure, the SMC’s gas gets an impulsive velocity kick of $\approx 30$ km s-1. Hence, ram pressure can potentially explain the lack of remnant rotation in gas (Figure 11), and the spatial offset between the SMC’s stellar density center and the gas LoS zero velocity curve (see Figure 1 right panel and Figure 12). The ram pressure exerted by the LMC’s ISM on the SMC during the SMC-LMC collision is much larger than the typical ram pressure values in galaxy CGMs (Emerick et al., 2016) and on the upper end of the ram pressure exerted by ICMs (Xie et al., 2025).